Assuming all volume measurements are made
at the same temperature and pressure, what
volume of hydrogen gas is needed to react
completely with 4.81 L of oxygen gas to produce water vapor?
Answer in units of L.
1
Expert's answer
2011-06-17T10:03:58-0400
The ballanced equation of producing water wapor is
2H2 + O2 = 2H2O
Since the volume of gas is proportional to the amount of moles, V(H2) = 2V(O2) = 4.81 * 2 = 9.62 L.